OpenStudy (anonymous):

How much should be deposited in an account paying 6% interest compounded semi-annually in order to have a balance of $13,500 after 3 years?

OpenStudy (stormfire1):

P = principal amount (the initial amount you borrow or deposit) r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year \[A = P(1+\frac{r}{n})^{nt}\] \[$13,500=P(1+\frac{.06}{2})^{2*3}\]Solve for P to get ~$11,306.04

OpenStudy (stormfire1):

Similar to your earlier posts, you just plug in all the know values and solve for the unknown one.